# From a Tiny Slip to a Giant Leap: An LLM-Based Simulation for Fake News Evolution

### v1

@@ 4. The prompt for the reasoning function is as:
5.The prompt for ”Official Statement” is as:
@@
Appendix B Human Evaluation
@@
Table 2. Correlation for LLM-based evaluations across FUSE-EVAL dimensions.
@@
The prompt used is as follows:
To validate the effectiveness of using LLMs for this task, we conduct a benchmarking study by comparing the LLM-generated evaluations with those from human judges. A group of human evaluators, knowledgeable in journalism and communication studies, independently assess a representative sample of the news content using the same scoring guidelines. We calculate the Pearson correlation coefficients between the scores assigned by the LLMs and the human evaluators for each dimension. The results, summarized in Table 2 , demonstrate that GPT-4o-mini exhibits a high correlation with human evaluations across all dimensions, outperforming gpt-3.5-turbo.
@@
Appendix C Alignment Between Simulated and Real-World Fake News
@@
Appendix D Various Topics and Simulation Results
Figure 6. The average deviation of news changes across different topics, social networks, dissemination role types, and traits.
In our experiments, we compared the evolution of fake news across five different topics: politics, science, finance, terrorism, and urban legends. As shown in Figure 6(a), political fake news spreads the fastest, with average deviation rapidly peaking within just four days and remaining at a high level. Fake news related to terrorism follows closely behind, showing similarly fast spread, likely due to the emotional intensity and urgency associated with such topics, which prompt individuals to quickly form beliefs and propagate the news widely. In contrast, financial news spreads at a slower pace, with deviation gradually accumulating over time. Although financial news is significant in terms of economic impact, individuals tend to engage in more rational thinking when encountering such news, leading to more stable growth in average deviation. Science-related fake news evolves the slowest, with average deviation consistently remaining low throughout the propagation process. These results is consistent with previous studies (Lazer et al., 2018a). This suggests that individuals are generally more cautious when dealing with scientific topics, often subjecting the information to more thorough verification.
@@ Appendix D Various Topics and Simulation Results
Appendix E Analysis of Experimental Costs
@@
Social networks in real life can generally be categorized into three types: high clustering networks, scale-free networks, and random networks, which correspond respectively to Figure 7 (a), (b), and (c).
Appendix F Simulation on Different Backbones
Figure 7. Different social network in our framework: (a) high clustering network (b) scale-free network (c) random network
Figure 8. Average Deviation changes with GPT-4 and GPT-4o-mini as the backbone under the terrorism topic, both of which demonstrate a deviation accumulation effect.
@@ Appendix F Simulation on Different Backbones
Appendix G Social Network
@@
where $\gamma$ is a parameter typically in the range 2 ¡ $\gamma$ 3. This distribution implies that most nodes have few connections, while a few hub nodes have a large number of connections. This heterogeneity in node connectivity is a hallmark of scale-free networks.
@@
$P(k)\approx\frac{\lambda^{k}e^{-\lambda}}{k!}$

### v2

@@ 4. The prompt for the reasoning function is as:
5.The prompt for “Official Statement” is as:
@@
Appendix C Implementation Details
Our simulation framework was developed using Python scripts, leveraging various libraries to model the agents and their environment effectively. The LLM used is gpt-4o-mini, accessed via OpenAI API calls. When creating the network structure, we used the Python library networkx to construct different social network structures. The simulation includes 40 agents, whose traits were based on the Big Five personality dimensions commonly used in psychology Barrick and Mount (1991). Each agent was assigned scores on these traits to introduce variability in behaviors and interactions within the simulation. For further details, please refer to our code at https://anonymous.4open.science/r/FUSE-7022/README.md.
Appendix D Human Evaluation
@@
To validate the effectiveness of using LLMs for this task, we conducted a benchmarking study by comparing LLM-generated evaluations with those from human judges. Three annotators (Ph.D. students in Computer Science and Technology and journalism studies) were recruited to independently assess a representative sample of 50 news items using the same scoring guidelines. We calculate the Pearson correlation coefficients between the scores assigned by the LLMs and the human evaluators for each dimension. The results, presented in Table 2, show that GPT-4o-mini achieves strong alignment with human evaluations across all dimensions, under a relatively high level of inter-annotator agreement (Fleiss’ $\kappa$ = 0.79) Mandrekar (2011), surpassing the performance of gpt-3.5-turbo.
The prompt used is as follows:
@@
Table 2: Correlation for LLM-based evaluations across FUSE-EVAL dimensions.
@@
Appendix E Alignment Between Simulated and Real-World Fake News
@@
Appendix F Various Topics and Simulation Results
Figure 6: The average deviation of news changes across different topics, social networks, dissemination role types, and traits.
In our experiments, we compared the evolution of fake news across five different topics: politics, science, finance, terrorism, and urban legends. As shown in Figure 6(a), political fake news spreads the fastest, with average deviation rapidly peaking within just four days and remaining at a high level. Fake news related to terrorism follows closely behind, showing similarly fast spread, likely due to the emotional intensity and urgency associated with such topics, which prompt individuals to quickly form beliefs and propagate the news widely. In contrast, financial news spreads at a slower pace, with deviation gradually accumulating over time. Although financial news is significant in terms of economic impact, individuals tend to engage in more rational thinking when encountering such news, leading to more stable growth in average deviation. Science-related fake news evolves the slowest, with average deviation consistently remaining low throughout the propagation process. These results is consistent with previous studies Lazer et al. (2018b). This suggests that individuals are generally more cautious when dealing with scientific topics, often subjecting the information to more thorough verification.
@@ Appendix F Various Topics and Simulation Results
Figure 7: Different social network in our framework: (a) high clustering network (b) scale-free network (c) random network.
Appendix G Analysis of Experimental Costs
@@
Social networks in real life can be categorized into three types: high clustering networks, scale-free networks, and random networks, which correspond respectively to Figure 7 (a), (b), and (c).
Appendix H Simulation on Different Backbones
Figure 8: Average Deviation changes with GPT-4 and GPT-4o-mini as the backbone under the terrorism topic, both of which demonstrate a deviation accumulation effect.
@@ Appendix H Simulation on Different Backbones
Appendix I Social Network
@@
where $\gamma$ is a parameter typically in the range 2 < $\gamma$ < 3. This distribution implies that most nodes have few connections, while a few hub nodes have a large number of connections. This heterogeneity in node connectivity is a hallmark of scale-free networks.
@@
$P(k)\approx\frac{\lambda^{k}e^{-\lambda}}{k!},$